OpenStudy (anonymous):

Solve for y. 2y – 16 = –4 A. 14 B. 6 C. –10 D. –18

OpenStudy (anonymous):

2y-16=-4 +16 +16 12=2y divide each side by 2 y=6

OpenStudy (anonymous):

Questions?

OpenStudy (anonymous):

I got It Thank you

OpenStudy (anonymous):

no prob!